OpenStudy (anonymous):

Solve for x. 2x2 - 4x = 0 0, -4 0, -2 0, 2 2, 4

myininaya (myininaya):

Factor the expression 2x^2-4x first

OpenStudy (anonymous):

factor out 2x, and you'll get 2x(x-2)=0 Now when you multiply two numbers, say A and B and together they are 0: A.B =0 Then what can you say for certain about at least one of the two?

OpenStudy (anonymous):

so is it b

myininaya (myininaya):

x-2=0 when x=?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Right, and what does x equal if 2x = 0?

OpenStudy (anonymous):

You have either 2x=0 or x-2=0... so two possible solutions